# DIRECT: Deep Active Learning under Imbalance and Label Noise

### v3

@@ 5.2 Experiments under Imbalance and Label Noise
We conduct novel sets of experiments under both class imbalance and label noise. Here, for both ResNet-18 and LabelBench experiments, we evaluate against all of the algorithms that performed well under the imbalance but noiseless setting above. For all of our experiments, we introduce a fixed percentage of label noise, where the given fraction of the examples’ labels are corrupted to a different class uniformly at random. For most of our experiements with $10\%$ label noise shown in Figure 5, we observe again that DIRECT consistently improves over all baselines including GALAXY. The results are consistent on ResNet-18 and LabelBench setups, and with different $B_{\text{parallel}}$ values, showing DIRECT’s robustness under label noise.
@@ 6 Conclusion and Future Work
In this paper, we conducted the first study of deep active learning under both class imbalance and label noise. We proposed an algorithm DIRECT that significantly and consistently outperforms existing literature. In this work, we also addressed the batch sampling problem of current SOTA algorithm, GALAXY (Zhang et al., 2022), by annotating multiple examples in parallel. Studying asynchronous labeling could be a natural extension of our work. A potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm. In addition, one can further batch the labeling process across different classes to further accommodate an even larger number of parallel annotators. Lastly, we refer the readers to Appendix F for our impact statement.
@@ Appendix C Time Complexity
In total, our experiments are conducted on NVIDIA 3090 ti GPUs. Each trial of the ResNet-18 experiment takes less than two hours while each trial of the LabelBench experiments takes roughly 12 hours.
@@ Appendix D Experiment Setup
ResNet-18 Experiments. ResNet-18 with passive training has been the standard evaluation in existing deep active literature [Ash et al., 2019, Zhang et al., 2022]. Our experiment setup utilizes the CIFAR-10, CIFAR-100 [Krizhevsky et al., 2009], SVHN [Netzer et al., 2011] and PathMNIST [Yang et al., 2021] image classification datasets. The original forms of these datasets are roughly balanced across 9, 10 or 100 classes. We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class. For example, given a balanced dataset above with $M$ classes. We generate an imbalanced dataset with $K$ classes ($K<M$) by the first $K-1$ classes from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$. Imbalance ratios are shown in Table 1.
For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer [Kingma and Ba, 2014]. The ResNet-18 model [He et al., 2016] is pretrained on ImageNet [Deng et al., 2009] from the PyTorch library. To address data imbalance, for all algorithms, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class. For experiments with label noise, we further add a $10\%$ label smoothing during training [Müller et al., 2019] for all algorithms.
LabelBench Experiments. Proposed by Zhang et al. [2024a], LabelBench evaluates active learning performance in a more comprehensive framework. Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model [Radford et al., 2021]. The framework also utilizes semi-supervised learning method FlexMatch [Zhang et al., 2021] to further leverage the unlabeled examples in the pool for training. We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW [Christie et al., 2018] and iWildcam [Beery et al., 2021]. Similar to the ResNet-18 experiments, for all algorithms, we use a $10\%$ label smoothing in the loss function to improve training under label noise. We did find FlexMatch to perform poorly under the combination of imbalance and label noise, so we used the passive training method for label noise experiments.

### v4

@@ 5.2 Experiments under Imbalance and Label Noise
We conduct novel sets of experiments under both class imbalance and label noise. Here, for both ResNet-18 and LabelBench experiments, we evaluate against all of the algorithms that performed well under the imbalance but noiseless setting above. For all of our experiments, we introduce a fixed percentage of label noise, where the given fraction of the examples’ labels are corrupted to a different class uniformly at random. For most of our experiments with $10\%$ label noise shown in Figure 5, we observe again that DIRECT consistently improves over all baselines including GALAXY. The results are consistent on ResNet-18 and LabelBench setups, and with different $B_{\text{parallel}}$ values, showing DIRECT’s robustness under label noise.
@@
5.3 Qualitative Analysis
The ImageNet-LT dataset is constructed with class frequencies that gradually decrease according to class index. In Figure 7, we organize classes into bins ordered from most frequent to least frequent. The results show that DIRECT significantly outperforms baseline algorithms on less frequent classes (indices 301-1000), which accounts for DIRECT’s superior overall performance despite modest accuracy reductions on more frequent classes. This finding aligns with our balancedness analysis, where DIRECT demonstrates improved labeling of samples from rare classes.
@@ 6 Conclusion and Future Work
In this paper, we conducted the first study of deep active learning under both class imbalance and label noise. We proposed an algorithm DIRECT that significantly and consistently outperforms existing literature. In this work, we also addressed the batch sampling problem of Zhang et al. (2022), by annotating multiple examples in parallel. Studying asynchronous labeling could be a natural extension of our work. A potential solution is to utilize an asynchronous variant of one-dimensional active learning algorithm. In addition, one can further batch the labeling process across different classes to further accommodate an even larger number of parallel annotators.
Impact Statement
@@ Appendix C Time Complexity
Below, we also provide a comprehensive list of computational complexity of different algorithms we consider in our implementation: As for data selection algorithms, let $K$ be the number of classes, $N$ be the pool size and $B_{\text{train}}$ be the batch size, $D$ be the penultimate layer embedding dimension and $T$ be the number of batches. Below, we detail the computation cost of data selection of each algorithm we consider.
  • DIRECT: $O(T(KN\log N+B_{train}N))$.
  • GALAXY: $O(T(KN\log N+B_{train}KN))$
  • BADGE: $O(TB_{train}N(K+D))$
  • Margin sampling/most likely positive/confidence sampling: $O(TKN)$
  • Coreset: $O(T^{2}B_{train}ND)$
  • SIMILAR: $O(TB_{train}ND)$
  • Cluster margin: $O(N^{2}\log N+TN(K+\log N))$
  • BASE: $O(TN(D+B_{train}))$
Overall, our experiments are conducted on NVIDIA 3090 ti GPUs. Each trial of the ResNet-18 experiment takes less than two hours while each trial of the LabelBench experiments takes roughly 12 hours.
@@ Appendix D Experiment Setup
ResNet-18 Experiments. ResNet-18 with passive training has been the standard evaluation in existing deep active literature (Ash et al., 2019; Zhang et al., 2022). Our experiment setup utilizes the CIFAR-10, CIFAR-100 (Krizhevsky et al., 2009), SVHN (Netzer et al., 2011) and PathMNIST (Yang et al., 2021) image classification datasets. The original forms of these datasets are roughly balanced across 9, 10 or 100 classes. We construct an extremely imbalanced dataset by grouping a large number of classes into one majority class. For example, given a balanced dataset above with $M$ classes. We generate an imbalanced dataset with $K$ classes ($K<M$) by the first $K-1$ classes from the original dataset and combining the rest of the classes $K,...,M$ into a single majority class $K$. Imbalance ratios are shown in Table 1. In addition, we also utilize the standard CIFAR-10LT and CIFAR-100LT variants in our experiments for noisy label setting.
For neural network training, we utilize the standard passive training on labeled examples with cross entropy loss and Adam optimizer (Kingma & Ba, 2014). The ResNet-18 model (He et al., 2016) is pretrained on ImageNet (Deng et al., 2009) from the PyTorch library. To address data imbalance, for all algorithms, we utilize a reweighted cross entropy loss by the inverse frequency of the number of labeled examples in each class. For experiments with label noise, we further add a $10\%$ label smoothing during training (Müller et al., 2019) for all algorithms.
LabelBench Experiments. Proposed by Zhang et al. (2024a), LabelBench evaluates active learning performance in a more comprehensive framework. Here, we fine-tune the large pretrained model from CLIP’s ViT-B32 model (Radford et al., 2021). The framework also utilizes semi-supervised learning method FlexMatch (Zhang et al., 2021) to further leverage the unlabeled examples in the pool for training. We conduct experiments on the two imbalanced datasets in LabelBench, with FMoW (Christie et al., 2018) and iWildcam (Beery et al., 2021). Similar to the ResNet-18 experiments, for all algorithms, we use a $10\%$ label smoothing in the loss function to improve training under label noise. We did find FlexMatch to perform poorly under the combination of imbalance and label noise, so we used the passive training method for label noise experiments.
@@ E.1 Noiseless Results under Imbalance
(c) ImageNet-LT Balanced Pool Accuracy
(d) iNaturalist Balanced Pool Accuracy
@@ E.2 Label Noise Results under Imbalance
(a) CIFAR-10LT, 10 classes, 15% label noise.
(b) CIFAR-100LT, 100 classes, 15% label noise.